# N-Agent Ad Hoc Teamwork

### v2

@@ 6.3 A Closer Look at POAM
Two key aspects of POAM are the use of a recurrent teammate modeling module, and the use of data from uncontrolled agents to train the critic (Q2). We perform further analysis of both aspects on the mpe-pp and 5v6 tasks; results on 5v6 may be found in Appendix A.4.
@@
In the NAHT training/evaluation setting, a new set of teammates is sampled at the beginning of each episode. Therefore, an important subtask for a competent NAHT agent is to rapidly model the distribution and type of teammates at the beginning of the episode, to enable the policy to exploit that knowledge as the episode progresses. This task is especially challenging in the presence of partial observability (a property of the SMAC tasks). To address the above challenges, POAM employs a recurrent encoder, which encodes the POAM agent’s history of observations and actions to an embedding vector, and a (non-recurrent) decoder network, which predicts the egocentric observations and actions for all teammates.
A natural question then, is whether the learned teammate encoding vectors actually improve over the course of an episode. Figures 4 and 9 depict the within-episode mean squared error (MSE) of the observation predicted by the decoder, and the within-episode probability of the action actually taken by the modelled teammates, according to the decoder’s modelled action distribution. Each curve corresponds to a different training checkpoint of a single run.
For both tasks, we observe that the average MSE decreases over the course of training, while the average probability of the taken actions increases. For later training checkpoints, we observe that the average MSE actually decreases within an episode, while the average probability increases. This reflects the increased confidence of the agent modelling module, as more data is observed about teammates. Thus, we conclude that POAM is able to cope with the challenges introduced by the sampled teammates and partial observability, to learn accurate teamate models.
@@
Figure 5: Learning curves of POAM and IPPO, where the value network is trained with and without uncontrolled agents’ data (UCD).
@@
6.4 Out of Distribution Generalization
Figure 6: Mean and standard error of test returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates. POAM shows improved generalization to OOD teammates.
The AHT literature commonly assumes that some subset of the uncontrolled teammates of interest may be interacted directly with during training [28, 6, 30]. Experiments in the previous sections were conducted under this assumption. This section conducts an analysis of how POAM performs when this assumption is relaxed—that is, when the teammates have not been seen previously (Q3). The out-of-distribution (OOD) teammate setting is of interest because in realistic scenarios, it may be challenging to enumerate all teammates likely to be encountered in the wild. Here, OOD teammates are generated by using MARL algorithms with different random seeds to generate teammates that have not been interacted with during training.
Figures 6 and 10 shows the mean and standard error of the test return achieved by POAM, compared to IPPO-NAHT, when the algorithm in question is paired with previously unseen seeds of IPPO, IQL, MAPPO, QMIX, and VDN In three out of five tasks (mpe, 5v6, 3s5z), POAM has a significantly higher return than IPPO-NAHT, while the remaining two tasks exhibit a smaller improvement. For all tasks, we observe that the std. error for POAM is lower than that of IPPO-NAHT, showing that POAM performs more consistently against a range of teammates and team configurations than IPPO-NAHT.
@@ 7 Related Works
This section summarizes related work in the areas most closely related to NAHT, namely, ad hoc teamwork, zero-shot coordination, agent modeling, and cooperative MARL.
@@
Agent modeling [2] helps us characterize other agents based on their actions. Such characterizations could attempt to infer important attributes of another agent’s decision making process such as their goals [12] or policies [28]. POAM and most existing AHT techniques rely on agent modeling to provide important teammate information for decision-making when collaborating with unknown teammates.
Cooperative MARL.
Cooperative Multi-Agent Reinforcement Learning (MARL) explores algorithms for training agent teams on fully cooperative tasks. Some existing methods focus on credit assignment and decentralized control [11, 32]. Other works in cooperative MARL also leverage parameter sharing and role assignment (e.g., [9, 42]) to decide an optimal division of labor between agents. However, these techniques assume control over all existing agents during training and evaluation. This limits these techniques’ effectiveness in settings with new teammates (NAHT problem) as shown in prior work [41, 15, 30].
@@ 8 Conclusion
This paper addresses a special case of the NAHT problem, with homogeneous and non-communicating agents. POAM, which employs full parameter sharing, may not perform well in settings with heterogeneous agents, or in settings that require highly differentiated roles. POAM also does not leverage centralized state information or allow communication between controlled agents. Incorporating this information might enable learning improved NAHT policies. Further, POAM’s actor update is purely on-policy, and therefore cannot leverage data generated by uncontrolled agents. Future work may consider off-policy learning methods as a solution strategy. In addition to the directions suggested by POAM’s limitations, algorithmic ideas from AHT, such as diversity-based teammate generation [24] and teammate-model-based planning methods [6], also suggest rich avenues for future work. Having introduced the NAHT problem in this work, we hope the community explores the many potential directions to design even better NAHT algorithms by considering advances in MARL, AHT and agent modeling.
@@ A.2 Further Discussion of the Motivating Example
We wish to analytically compute the global maxima of this expression with respect to $p$, by considering the boundary points $p\in{0,1}$ and the zeros of the first derivative.
@@ A.3.1 Team Sampling Procedure Used in Experiments
Recall that $U$ and $C$ denote the sets of uncontrolled and controlled agent policies, respectively, while $M$ denotes the total team size, and $N$ the number of agents sampled from $C$. The experiments in the following consider an $X_{train}$ consisting of sampling $N$ uniformly from $\{1,\cdots,M-1\}$, sampling $N$ agents from $C$ and $M-N$ agents from $U$ in a uniform fashion (Figure 7). The sampling procedure takes place at the beginning of each episode to select a team, which is then deployed in the environment. Data generated by the deployed team (e.g. joint observations, joint actions, and rewards) are returned to the learning algorithm. See Appendix A.3.3 for further details on the evaluation procedure.
@@ A.3.2 Experimental Domains
The MPE environment [23] (released under the MIT License) is a setting where particle agents interact within a bounded 2d plane, equipped with a discrete action space and continuous observation space. The observation space is 16-dimensional, and contains the agents’ own position and velocity, relative positions/velocities of all other agents, landmarks, and prey. The action space consists of 5 discrete actions, corresponding to the four cardinal movement directions and a no-op action.
The predator-prey task (mpe-pp) is a custom task implemented by the authors of this paper within the fork of the MPE environment released by Papoudakis et al. [29] (MIT License). In this task, three predators must cooperatively pursue a pre-trained prey agent. The team receives a reward of 1 if two or more predators collide with the prey at a single time step, and no reward if only a single agent collides with the prey. A shaping reward consisting of 0.01 times the $\ell_{2}$ distance between each agents in the team and the prey is provided as well. The maximum episode length is 100 time steps.
@@
SMAC [33] (released under the MIT License) features a set of cooperative tasks, where a team of allied agents must defeat a team of enemy agents controlled by the game server. It is a partially observable domain with a continuous state space and discrete action space. For each agent, the observation space is continuous, and consists of features about itself, enemy, and allied agents within some radius. The action space is discrete, and allows an agent to choose an enemy to attack, a direction to move in, or to not perform any action. As the number of allies and enemies varies between tasks, the dimensionality of the observation and action spaces is particular to each task. The maximum number of time steps per episode is also task specific, although early termination occurs if all enemies are defeated. At each time step, the team receives a shaped reward corresponding to the damage dealt, and bonuses of 10 and 200 points for killing an enemy and winning the scenario by defeating all enemies. The reward is scaled such that the maximal return achievable in each task is 20.
@@
Figure 8: Evolution of the within-episode mean squared error (left) and probability of actions that were actually taken by modelled teammates (right), computed by the POAM agent over the course of training, on the 5v6 task. The agent modeling performance improves over the course of an episode, as more data about teammate behavior is observed.
Figure 10: Mean and standard error of test returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates. POAM shows improved generalization to OOD teammates, as compared to IPPO-NAHT, across all tasks.
@@ A.3.3 Evaluation Details
The experiments in the main paper consider a particular team sampling procedure called $X_{train}$, which uniformly samples $N$ and uniformly samples agents from the sets $U$ and $C$. While the learning curves reported in the main paper are computed directly with $X_{train}$, we use an iterative and deterministic evaluation procedure for the results presented in Section 6.4 and in the Appendix. The purpose of the additional evaluation procedure presented here is to ensure that we thoroughly evaluate the trained NAHT agents for all possible values of $N$ and all possible teammates.
Given a set of uncontrolled agents $U$ and a set of agents $C$ generated by some NAHT method, we compute the following $M-N$ cross-play scores. Let $N$ be the number of agents sampled from $C$, such that $N<M$. For $N\in\{1,\cdots,M-1\}$, construct the joint policy $\bm{\pi}^{(N)}$ by selecting $N$ agents uniformly from $C$ and $M-N$ agents from $U$. Evaluate the resultant team on the task for $E$ episodes. This results in $M-1$ episode returns, on which we may compute summary statistics. If agents in $U$ and $C$ are generated by algorithm $A$ and $B$, respectively, then we call the thus computed average return, the $M-N$ cross-play score.
More concretely: each algorithm (VDN, QMIX, QMIX-NS, IQL, IQL-NS) is run $k$ times with different seeds, to generate $k$ teams of teammates to act as uncontrolled agents. For each pair of algorithms, we sample a subset of the possible seed pairs, and evaluate the teams that result from merging said seed pairs. Note that each seed actually represents a team of agents. For example, if VDN and QMIX have seeds 1 2, and 3, the cross-play evaluation might consider the seed pairings (VDN 1, QMIX 2), (VDN 2, QMIX 3), (VDN 3, QMIX 1). Given a pair of seeds (e.g. (VDN 1), (QMIX 2)), the $M-N$ cross-play score is computed as the average return generated by sweeping $N\in\{1,\cdots,M-1\}$ and evaluating the merged team that consists of selecting $N$ agents from the first team, and $M-N$ agents for $E$ episodes. In our experiments, $E=128$, and $k=3$.
Means and standard errors are computed over all seed pairings considered, as we treat each $M-N$ evaluation as an independent trial. For most experiments, three seed pairs are considered, where seeds are paired to ensure that all seeds participate in at least one evaluation. This ensures that the computational cost of the evaluation remains linear in $k$, rather than quadratic. However, note that for the OOD experiments presented in Section 6.4, we consider all possible seed pairings for the most comprehensive evaluation.
@@
For algorithm pairs that are the same, seeds are paired with themselves. For example, the seed pairs for VDN versus VDN would be $(1,1),(2,2),(3,3)$. The reason for this is that we are interested in the performance of agents when paired with known teammates. This is the model self-play score, in contrast to the algorithm self-play score [3].
@@ A.3.4 Algorithm Implementation
The experiments in this paper use algorithm implementations from the ePyMARL codebase [29] (released under the Apache License). The value-based methods are used without modification (i.e. IQL, VDN, QMIX), but we implement our own version of policy gradient methods (IPPO, MAPPO), based on the implementation of [43].
All methods employ recurrent actors and critics, with full parameter sharing, i.e. all agents are controlled by the same policy, where the agent id is input to the policy/critic networks, to allow behavioral differentiation between agents. All methods employ the Adam optimizer to optimize all networks involved. For policy gradient methods, the policy architecture is two fully connected layers, followed by an RNN (GRU) layer, followed by an output layer. Each layer has 64 neurons with ReLU activation units, and employs layer normalization. The critic architecture is the same as the policy architecture. The value-based networks employ the same architecture, except that there is only a single fully connected layer before the RNN layers, and layer normalization is not used, following the ePyMARL implementation. Please consult the code base for full implementation details.
Hyperparameters.
For the value-based methods, default hyperparameters are used. We tune the hyperparameters of the policy gradient methods on the 5v6 task, and apply those parameters directly to the remaining SMAC tasks. The hyperparameters considered for policy gradient algorithms are given in Table 1. POAM adopts the same hyperparameters as IPPO where applicable. We also tuned additional hyperparameters specific to POAM (see Table 2).
@@
A.4 Supplemental Results
A.4.1 Validating the Existence of Coordination Conventions

### v3

@@ 6.4 A Closer Look at POAM
Two key aspects of POAM are the teammate modeling module, and the use of data from uncontrolled agents to train the critic (Q3). We study the impact of both aspects on POAM’s sample efficiency, focusing on the mpe-pp and 5v6 tasks. Results on 5v6 may be found in Appendix A.4.
@@
In the NAHT training/evaluation setting, a new set of teammates is sampled at the beginning of each episode. Therefore, an important subtask for a competent NAHT agent is to rapidly model the distribution and type of teammates at the beginning of the episode, to enable the policy to exploit that knowledge as the episode progresses. This task is especially challenging in the presence of partial observability (a property of the SMAC tasks). To address the above challenges, POAM employs a recurrent encoder, which encodes the POAM agent’s history of observations and actions to an embedding vector, and a (non-recurrent) decoder network, which predicts the egocentric observations and action distribution for all teammates.
Figure 5: Evolution of a POAM agent’s within-episode mean squared error (left) and within-episode probability of actions of modeled teammates (right), over the course of training on mpe-pp.
A natural question then, is whether the learned teammate embedding vectors actually improve over the course of an episode. Figures 5 and 9 depict the within-episode mean squared error (MSE) of the observation predicted by the decoder, and the within-episode probability of the action actually taken by the modeled teammates, according to the decoder’s modeled action distribution. Each curve corresponds to a different training checkpoint of a single run.
For both tasks, we observe that the average MSE decreases over the course of training, while the average probability of the taken actions increases. For later training checkpoints, we observe that the average MSE actually decreases within an episode, while the average probability increases, reflecting the increased confidence of the agent modelling module as more data is observed about teammates. Thus, we conclude that POAM is able to cope with the challenges introduced by the sampled teammates and partial observability, to learn accurate teammate models.
@@
Figure 6: Learning curves of POAM and IPPO-NAHT, where the value network is trained w/w.o. uncontrolled agents’ data (UCD).
@@ 7 Related Works
This section summarizes literature in areas most closely related to NAHT, namely, ad hoc teamwork, zero-shot coordination, evaluation of cooperative capabilities, agent modeling, and CMARL.
@@
Agent modeling enables agents to characterize other agents based on their actions [2]. Such characterizations could attempt to directly infer modeled agents’ actions, goals [13], or policies [29]. The modeled attributes have been used in cooperative, competitive, and general sum settings to inform update rules [11, 36] or to directly inform decision making [29]. POAM relies on agent modeling to provide important teammate information for decision-making when collaborating with unknown teammates.
Cooperative MARL (CMARL).
CMARL explores algorithms for training agent teams on fully cooperative tasks. Some existing methods focus on credit assignment and decentralized control [12, 33]. Other works in CMARL also leverage parameter sharing and role assignment (e.g., [9, 44]) to decide an optimal division of labor between agents. However, these techniques assume control over all existing agents during training and evaluation, which limits their effectiveness in settings with unseen or uncontrolled teammates, as shown in prior work [43, 16, 31].
@@ 8 Conclusion
This paper addresses a special case of the NAHT problem, with homogeneous and non-communicating agents. POAM, which employs full parameter sharing, may not perform well in settings with heterogeneous agents, or in settings that require highly differentiated roles. POAM also does not leverage centralized state information or allow communication between controlled agents. Incorporating this information might enable learning improved NAHT policies. Further, POAM’s actor update is purely on-policy, and therefore cannot leverage data generated by uncontrolled agents. Future work might consider employing off-policy methods to exploit the uncontrolled agent data. In addition to the directions suggested by POAM’s limitations, algorithmic ideas from AHT, such as diversity-based teammate generation [25] and teammate-model-based planning methods [6], also suggest rich avenues for future work. Having introduced the NAHT problem in this work, we hope the community explores the many potential directions to design even better NAHT algorithms by considering advances in MARL, AHT, and agent modeling.
Acknowledgments and Disclosure of Funding
This work has taken place in the Learning Agents Research Group (LARG) at UT Austin. LARG research is supported in part by NSF (FAIN-2019844, NRT-2125858), ONR (N00014-18-2243), ARO (W911NF-23-2-0004, W911NF-17-2-0181), Lockheed Martin, and UT Austin’s Good Systems grand challenge. Peter Stone serves as the Executive Director of Sony AI America and receives financial compensation for this work. The terms of this arrangement have been reviewed and approved by the University of Texas at Austin in accordance with its policy on objectivity in research.
@@ A.2 Further Discussion of the Motivating Example
We wish to analytically compute the global maxima of this expression with respect to $p$, by considering the boundary points $p\in\{0,1\}$ and the zeros of the first derivative.
@@ A.3.1 Team Sampling Procedure Used in Experiments
Recall that $U$ and $C$ denote the sets of uncontrolled and controlled agent policies, respectively, while $M$ denotes the team size for the task, and $N$ the number of agents sampled from $C$. The experiments in the following consider an $X_{train}$ consisting of sampling $N$ uniformly from $\{1,\cdots,M-1\}$, sampling $N$ agents from $C$ and $M-N$ agents from $U$ in a uniform fashion (Figure 7). The sampling procedure takes place at the beginning of each episode to select a team, which is then deployed in the environment. Data generated by the deployed team (e.g. joint observations, joint actions, and rewards) are returned to the learning algorithm. See Appendix A.3.3 for further details on the evaluation procedure.
@@ A.3.2 Experimental Domains
The MPE environment [24] (released under the MIT License) is a setting where particle agents interact within a bounded 2D plane, equipped with a discrete action space and continuous observation space. The observation space is 16-dimensional, and contains the agents’ own position and velocity, relative positions/velocities of all other agents, landmarks, and prey. The action space consists of five discrete actions, corresponding to the four cardinal movement directions and a no-op action. The observation range is normalized to $[-1,1]$, while the discrete action space is one-hot encoded.
The predator-prey task (mpe-pp) is a custom task implemented by the authors of this paper within the fork of the MPE environment released by Papoudakis et al. [30] (MIT License). In this task, three predators must cooperatively pursue a pre-trained prey agent. We use the pre-trained prey policy provided by the ePymarl MARL framework. The prey policy was originally trained by Papoudakis et al. [30], by using the MADDPG MARL algorithm to train both predator and prey agents for 25M steps, and generally attempts to escape approaching predators. The team receives a reward of 1 if two or more predators collide with the prey at a single time step, and no reward if only a single agent collides with the prey. A shaping reward consisting of 0.01 times the $\ell_{2}$ distance between each agents in the team and the prey is provided as well. Since the prey policy is pre-trained and fixed, note that our predator-prey task is a fully cooperative task from the perspective of the predator (learning) agents. The maximum episode length is 100 time steps.
@@
SMAC [34] (released under the MIT License) features a set of cooperative tasks, where a team of allied agents must defeat a team of enemy agents controlled by the game server. It is a partially observable domain with a continuous state space and discrete action space. For each agent, the observation space is continuous, and consists of features about itself, enemy, and allied agents within some radius. The action space is discrete, and allows an agent to choose an enemy to attack, a direction to move in, or to not perform any action. As the number of allies and enemies varies between tasks, the dimensionality of the observation and action spaces is particular to each task. The observation space is normalized to be between $[-1,1]$, while the action space is one-hot encoded. The maximum number of time steps per episode is also task specific, although early termination occurs if all enemies are defeated. At each time step, the team receives a shaped reward corresponding to the damage dealt, and bonuses of 10 and 200 points for killing an enemy and winning the scenario by defeating all enemies. The reward is scaled such that the maximal return achievable in each task is 20.
@@ A.3.3 Evaluation Details
This section provides details on how mixed teams are evaluated in the NAHT setting, and how cross-play scores, and self-play scores, and uncertainty measures are computed.
$M-N$ score.
Given a set of controlled agents $C$, and a set of uncontrolled agents $U$, the goal of the $M-N$ score is to quantify the performance when these two teams must cooperate within the NAHT scenario. The $M-N$ score is computed in a deterministic and exhaustive fashion, by iterating over all possible values of $N$. Let $N$ be the number of agents sampled from set $C$, such that $N<M$. For $N\in\{1,\cdots,M-1\}$, construct the joint policy $\bm{\pi}^{(M)}$ by selecting $N$ agents uniformly from $C$ and $M-N$ agents from $U$. Evaluate the resultant team on the task for $E$ episodes. This results in $(M-1)*E$ episode returns, which is averaged to form the $M-N$ score.
Cross-play scores.
The cross-play scores reported in this paper are the average returns of teams generated by algorithm $A$, when coordinating with those generated by $B$. To compute the cross-play score, we first train multiple teams (by varying the seed) via both algorithms $A$ and $B$. Next, the $M-N$ score is computed for random pairings of teams (seeds) from $A$ and $B$, following the procedure specified above. Summary statistics can be computed over the set of all such NAHT returns.
For example: each algorithm (VDN, QMIX, IQL, MAPPO, IPPO) is run $k$ times with different seeds, to generate $k$ teams of agents that may act as uncontrolled teammates. For each pair of algorithms, we sample a subset of the possible seed pairs, and evaluate the teams that result from merging said seed pairs. If VDN and QMIX have seeds 1 2, and 3, the cross-play evaluation might consider the seed pairings (VDN 1, QMIX 2), (VDN 2, QMIX 3), (VDN 3, QMIX 1). Given a pair of seeds (e.g. (VDN 1), (QMIX 2)), the $M-N$ cross-play score is computed as the average return generated by sweeping $N\in\{1,\cdots,M-1\}$ and evaluating the merged team that consists of selecting $N$ agents from the first team, and $M-N$ agents for $E$ episodes. In our experiments, $E=128$, and $k=5$.
@@
The self-play scores reported in this paper are model self-play score, rather than algorithm self-play score [3]. The reason for this is that we are interested in the performance of agents when paired with known teammates.
Measuring uncertainty
Means and 95% confidence intervals are computed over all team/seed pairings considered for any given scores, as we treat each $M-N$ evaluation as an independent trial. For most experiments, five seed pairs are considered, where seeds are paired to ensure that all seeds participate in at least one evaluation. This ensures that the computational cost of the evaluation remains linear in $N$, rather than quadratic. However, note that for the OOD experiments presented in Section 6.3, we consider all possible seed pairings for the most comprehensive evaluation.
@@ A.3.4 Algorithm Implementation
The experiments in this paper use algorithm implementations from the ePyMARL codebase [30] (released under the Apache License). The value-based methods are used without modification (i.e. IQL, VDN, QMIX), but we implement our version of policy gradient methods (IPPO, MAPPO), based on the implementation of Yu et al. [45].
Parameter Sharing.
All methods employ recurrent actors and critics, with full parameter sharing, i.e. all agents are controlled by the same policy, where the agent id is input to the actor and critic networks, to allow behavioral differentiation between agents. For POAM, which also maintains encoder-decoder networks for agent modeling, parameter sharing is used for the encoder and decoder networks, to improve training sample efficiency. Further, POAM’s decoder, which predicts observations and actions for all $-i$ agents, employs parameter sharing across agent predictions, to prevent the target dimensionality from scaling with the number of teammates.
Optimizer and Neural Architecture.
The Adam optimizer is applied for all networks involved. For policy gradient methods, the policy architecture is two fully connected layers, followed by an RNN (GRU) layer, followed by an output layer. Each layer has 64 neurons with ReLU activation units, and employs layer normalization. The critic architecture is the same as the policy architecture. The value-based methods employ the same architecture, except that there is only a single fully connected layer before the RNN layers, and layer normalization is not used, following the ePyMARL implementation. Please consult the codebase for full implementation details.
@@
Hyperparameters.
For the value-based methods, default hyperparameters are used. We tune the hyperparameters of the policy gradient methods on the 5v6 task, and apply those parameters directly to the remaining SMAC tasks. The hyperparameters considered for policy gradient algorithms are given in Table 1. POAM adopts the same hyperparameters as IPPO where applicable. We also tuned additional hyperparameters specific to POAM (Table 2).
@@
A.4 Supplemental Figures
This section contains additional figures referenced by the main paper. Please see Section 6 for the corresponding analysis and discussion.
Figure 8: Evolution of the within-episode mean squared error (left) and probability of actions that were actually taken by modeled teammates (right), computed by the POAM agent over the course of training, on the 5v6 task. The agent modeling performance improves over the course of an episode, as more data about teammate behavior is observed.
Figure 10: Returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates. POAM shows improved generalization to OOD teammates, as compared to IPPO-NAHT, across all StarCraft tasks. For each type of teammate, the performance of IPPO-NAHT/POAM against the exact teammate seen during training is shown as the in-distribution baseline.
A.5 Supplemental Analysis
The following subsections contains secondary analysis and results, intended to support the primary analysis of the main paper. For all results, the mean and 95% confidence interval over five trials is reported. All reported returns are test returns.
A.5.1 The Need for Dedicated NAHT Algorithms - Empirical Evidence
Section 4 argues theoretically that with a population of uncontrolled teammates that select one with probability 1/3 and zero otherwise, an optimal policy learned in the AHT setting would not be optimal in the NAHT setting. Here, we present experiments on the three agent bit matrix game that empirically verify the theory. Let $N$ denote the number of controlled agents. The episode length is 25, and the observation for each agent consists of the agent index and the joint action at the previous time step. The team reward at each time step is $3*\mathbbm{1}_{\sum_{i}b_{i}=1}$.
The optimal expected return in the $N=1$ (AHT) setting is 33.333 (derived from Lemma A.3), and in the $N=2$ (NAHT) setting is 50.0 (derived from Lemma A.4). The methods compared are POAM-AHT and POAM. Table 3 shows the expected returns achieved by POAM and POAM-AHT on the $N=1$ and $N=2$ scenarios. Since any policy is optimal in the $N=1$ case, as expected, both methods achieve near-optimal returns. In the $N=2$ case, as expected, POAM outperforms POAM-AHT by a large margin, achieving a near-optimal return of $48.858\pm 1.092$.
N=1 (AHT) N=2 (NAHT)
POAM-AHT 33.441 $\pm$ 0.511 22.5 $\pm$ 8.250
POAM 33.483 $\pm$ 0.485 48.858 $\pm$ 1.092
Table 3: Returns of POAM-AHT versus POAM on the three agent bit matrix game, in the N=1 (AHT) and N=2 (NAHT) setting. POAM and POAM-AHT both achieve the optimal return for the $N=1$ case, but POAM has a much higher return on the $N=2$ case.
A.5.2 Validating the Existence of Coordination Conventions